OpenStudy (anonymous):

Select the ordered pair from the choices below that is a solution to the following system of equations: 2x = y + 11 3y + 7 = -(1/2)x A. (-2, -15) B. (10, 9) C. (4, -3) D. (-1, -13)

OpenStudy (kropot72):

If you multiply the second equation by a number you can make the term in x have the same value as the term in x in the first equation, but with a negative sign. The like terms in the equations can then be added and the terms in x will be eliminated. What number can be used as a multiplier?

OpenStudy (kropot72):

Do you understand the above explanation?

OpenStudy (anonymous):

noo

jimthompson5910 (jim_thompson5910):

Start with 2x = y + 11 and solve for y 2x = y + 11 2x-11 = y + 11-11 2x-11 = y y = 2x-11 ---------------------- Now because y = 2x-11, we can replace 'y' in the second equation with 2x-11 (since they are the same thing) 3y + 7 = -(1/2)x 3(2x-11) + 7 = -(1/2)x Now solve for x 3(2x-11) + 7 = -(1/2)x 6x-33 + 7 = -(1/2)x 6x-26 = -(1/2)x 2*6x-2*26 = -2*(1/2)x ... Note: Multiply EVERY term by the LCD 2 to clear out the fraction 12x - 52 = -x -52 = -x-12x -52 = -13x -52/(-13) = x 4 = x x = 4 ---------------------------------------------- Now that we know that x = 4, we can use this to find y. Plug it into y = 2x-11 to find y. y = 2x-11 y = 2(4) - 11 y = 8-11 y = -3 So the solutions are x = 4 and y = -3 So the solution as an ordered pair is (4,-3)

OpenStudy (anonymous):

thanks that helped

OpenStudy (kropot72):

I am confused. Should we be presenting model solutions thru to an answer or should we go step by step with the questioner?